With 50% margin, what is his/her return over a twelve-month period if the stock depreciates from $100 to $50 (assuming no annual interest rate on margin)?  What is the return assuming no margin?
Answer #1

With 50% margin

Initial investment = $100*50% = $50

Return = (Price after 1 years- Initial investment)/ Initial investment

= (50-50)/50

= 0

With No Margin

Initial investment = $100

Return = (Price after 1 years- Initial investment)/ Initial investment

=(50-100)/100

= -50%